OpenStudy (anonymous):

Flying against the wind, an airplane travels 4620km in 6 hours. Flying with the wind, the same plane travels 3750km in 3 hours. What is the rate of the plane in still air and what is the rate of the wind?

OpenStudy (anonymous):

The unit rate or just the rate?

OpenStudy (amistre64):

velocity is distance divided by time .... setup the 2 equations, in 2 unknowns

OpenStudy (ddcamp):

Using a variable for the plane's speed and another for the windspeed, can you come up with equations for the total velocity when the plane is (a) travelling with the wind, and (b) travelling against the wind?

OpenStudy (anonymous):

(P+w)3 = 3750 (p-w)6 = 4620 this correct?

OpenStudy (radar):

Looks like a good start.

OpenStudy (radar):

Now set up your system and decide which variable you will eliminate w or p. Whatever you choose, you will still need to solve the other one as the problem is asking for both the wind and aircraft speed.